# On a fast consistent selection of nested models with possibly unnormalised probability densities

### v1

@@ 1 Introduction
At present, there are essentially two main approaches to overcome the challenge. One approach is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo. See, for example, (Baker, 2022) and (MacKay, 2003; Congdon, 2006). While simple and intuitive, this approach is computationally intensive, especially for random vectors in high-dimensional settings. Another approach is to avoid the calculation of the normalizing constant altogether. For this, the score matching method (Hyvärinen, 2005) provides significant potential through efficient computation, as it does not depend on the normalizing constant. This method has been applied to Bayesian model selection (Dawid and Musio, 2015; Shao et al., 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution. In this approach, consistent selection is only proved for non-nested models ((Shao et al., 2019), p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear. Our approach is different. We allow the data to be sampled directly from an unnormalized PDF.
@@ 2.1 Fisher Divergence and Gentropy
$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0.$
@@
Next, within their new entropy structure, Cheng and Tong (2024) define, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
@@
Denote $x$ and $y$ as two random vectors with joint PDF $p(x,y)$. Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$. Tables 1 and 2 summarize the comparison between Shannon entropy and G-entropy across different definitions and properties.
@@
Entropy $H(x)=H(p)=E_{p}[Loss(p(x))]$ $H_{G}(x)=H_{G}(p)=E_{p}||\nabla_{x}\log p(x)||^{2}$
Cross entropy $H(p,q)=E_{p}[Loss(q(x))]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
@@ 2.2 GIC Estimate
Following equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$. When $p$ is the true data PDF and $q$ the model PDF, $H_{G}(p,q)$ inspires the maximum GIC estimate (MGICE).
@@ 2.3 GIC for Model Selection
Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $nGIC$ as follows:
@@
where $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$, $GIC_{n}(M_{k}(\hat{\theta}_{k}))=nGIC(M_{k}(\hat{\theta}_{k}))$, and
@@ 3.2 Consistency of T-GIC
Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenious norm. Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
@@ 4.1 $N\times t$ PDF
We conduct 100 replications to evaluate the performance of the MGICE, with sample size $n=1000,3000,\text{and }5000$, where the true parameter values are set to $(\mu^{*},s^{*},\alpha^{*},k^{*})=(0.3,0.5,0.5,1.5)$. Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution. For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm to jointly optimize all parameters. Similar to Baker (2022), for all experiments involving the $N\times t$ PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g. $\alpha=0.25$ or 1, and $k=1$ or 2. Table 4 presents the average MGICE and its SD, showing good overall consistency of MGICE, although there is still room for improvement for the parameters $s^{*},\alpha^{*},$ and $k^{*}$.
@@ 4.2 AR Model with $N\times t$ Noise
We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $N=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c,s,\alpha,k)$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$. In each replication, the first 200 data are discarded to ensure stationarity.
@@ 4.3 Polynomial Regression Model with $N\times t$ Distributed Errors
We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $n=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c,s,\alpha,k)$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
@@ 4.4 A Bivariate Model with a Von Mises PDF
We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using T-GIC1 and T-GIC2, with sample size $n=300,500,\text{and }1000$. The true parameter dimension is set to 5, with true parameter $\theta^{*}=(2.0,1.0,1.5,2.5,3.0)^{T}$. Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $[0,2\pi)\times[0,2\pi)$. For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm to jointly optimize all parameters, with all parameters initialized to zero.
@@ 5.1 Finance Data
where $\varepsilon_{t}\sim N(0,\sigma^{2})$. Tables 10 and 11 summarize the results. Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model. However, under $N\times t$ noise, T-GIC1 selects a sixth-order AR model, while T-GIC2 selects a second-order AR model, both yielding estimates of $\alpha$ and $k$ similar to Baker (2022). Since AR models offer a convenient framework for prediction, Table 12 presents a comparison of out-of-sample performance between the two fitted models, either chosen by AIC vs BIC or by T-GIC1 vs T-GIC2. In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling k-step-ahead forecast. It is interesting to observe, for T-GIC1 vs T-GIC2, that the fitted AR(6) model performs better in forecasting for the short term than the fitted AR(2) model but worse for the longer term. However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly.
@@
Table 12: MSE and the ratio in the rolling k-step-ahead forecast on AR models.
$k$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
@@ A.2 Proof of Theorem 1
First, we observe that for any model $M$ satisfying Assumptions 6-7 and assuming the validity of the law of large numbers, then it follows from routine analysis that $GIC(M(\theta))=P_{n}\{W(x,p_{M(\theta)})\}$ converges uniformly in probability to its population version $GIC_{\infty}(M(\theta))=P\{W(x,p_{M(\theta)})\}$. It follow from Assumption 7 that $GIC_{\infty}(M_{k}(\theta_{k}))$ is a Lipschitz-continuous, as a function of the parameter $\theta_{k}$, hence it attains its maximum value, denoted by $\mathcal{M}_{k}$, owing to the compact parameter space assumption (aka Assumption 6). Suppose $k_{0}$ is the smallest $k$ such that $M_{k}$ contains the true model. Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise. Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{F}(p^{*})>0$ where $p^{*}$ is the true population pdf. Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum T-GIC model selection criterion will not select any $k<k_{0}$, in probability.
Henceforth, consider the case that $k\geq k_{0}$. Since $\mathcal{M}_{k_{0}}>0$, $GIC(k)$ is positive, in probability, i.e., $GIC(k)>0$ holds with probability approaching 1 as sample size increases without bound. For ease of exposition, we shall assume that $GIC(k)$ is positive. Let $k>k_{0}$ be fixed. Consider the increment $D=\log\{T\text{-}GIC(k)\}-\log\{T\text{-}GIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$. By Proposition 1, $n\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution. Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to-\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$. This completes the proof of the consistency of the proposed T-GIC model selection criterion.

### v2

@@ 1 Introduction
At present, there are essentially two main approaches to overcome the challenge. One approach is to first estimate the intractable normalizing constant by numerical approximations or Markov Chain Monte Carlo. See, for example, (Baker, 2022) and (MacKay, 2003; Congdon, 2006). While simple and intuitive, this approach is computationally intensive, especially for random vectors in high-dimensional settings. Another approach is to avoid the calculation of the normalizing constant altogether. For this, the score matching method (Hyvärinen, 2005) provides significant potential through efficient computation, as it does not depend on the normalizing constant. This method has been applied to Bayesian model selection (Dawid and Musio, 2015; Shao et al., 2019) that focuses mainly on the issue of improper priors while the data are still sampled from a normalized distribution. In this approach, consistent selection is only proved for non-nested models (Shao et al., 2019, p.1826), with the situation for nested models, such as polynomial regression models or autoregressive (AR) models, remaining unclear. Our approach is different. We allow the data to be sampled directly from an unnormalized PDF.
@@ 2.1 Fisher Divergence and Gentropy
$p(x_{1},\cdots,x_{i-1},-\infty,x_{i+1},\cdots,x_{d})\equiv 0\text{ and }p(x_{1},\cdots,x_{i-1},+\infty,x_{i+1},\cdots,x_{d})\equiv 0,$
where, e.g., $p(x_{1},\cdots,x_{i-1},\infty,x_{i+1},\cdots,x_{d})$ denotes $\lim_{x_{i}\to\infty}p(x_{1},\cdots,x_{d})$.
@@
Next, within their new entropy structure, Cheng and Tong (2024) defined, among other notions, Gentropy $H_{G}(p)$ and G-cross-entropy $H_{G}(p,q)$ as
@@
Denote $x$ and $y$ as two random vectors with joint PDF $p(x,y)$. Let $\theta$ be the unknown $h$-dimensional parameter vector for $p(x)$, $h\geq 1$. Tables 1 and 2 summarize a comparison between the Shannon entropy and the G-entropy (in terms of definitions and properties). The tables are based on the results developed by Cheng and Tong (2024).
@@
Entropy $H(x)=H(p)=E_{p}[-\log p(x)]$ $H_{G}(x)=H_{G}(p)=E_{p}[W(x,p)]$
Cross entropy $H(p,q)=E_{p}[-\log q(x)]$ $H_{G}(p,q)=E_{p}[W(x,q)]$
@@ 2.2 GIC Estimate
By equation (2.1), given PDF $p$, minimizing $D_{F}(p||q)$ with respect to $q$ is equivalent to maximizing the G-cross-entropy $H_{G}(p,q)$. When $p$ is the true data PDF and $q$ the model PDF, $H_{G}(p,q)$ inspires the maximum GIC estimate (MGICE).
@@ 2.3 GIC for Model Selection
Consider a collection of candidate parametric models $M_{1},\cdots,M_{K}$, denoted as $M_{k}(\theta_{k}),k=1,\cdots,K$, with $\theta_{k}$ ($h_{k}\geq 1$) being an unknown $h_{k}$-dimensional parameter vector in parameter space $\Theta.$ Similar to AIC, an unbiased $GIC_{c}$ for model selection (Cheng and Tong 2024) is derived under mild regularity conditions by correcting the bias, $B$, in $n\times GIC$ as follows:
@@
where $\hat{\theta}_{k}$ is the MGICE of $\theta_{k}$, $GIC_{n}(M_{k}(\hat{\theta}_{k}))=n\times GIC(M_{k}(\hat{\theta}_{k}))$, and
@@ 3.2 Consistency of T-GIC
Both $\nabla_{x}^{2}\log p_{M(\theta)}(x)$ and $[\nabla_{x}\log p_{M(\theta)}(x)][\nabla_{x}\log p_{M(\theta)}(x)]^{T}$ are Lipschitz continuous in respect of Frobenius norm. Specifically, $\forall\theta_{1},\theta_{2}\in\Theta$,
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. If $K=k_{0}$, under Assumptions 1-8, then $T\textbf{-}GIC1=\exp\{-2\#(M_{k})/n\}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
@@
Let $M_{1},M_{2},\cdots,M_{K}$ be a finite sequence of strictly nested models, with $k_{0}$ being the smallest $1\leq k\leq K$ such that $M_{k}$ contains the true model. Under Assumptions 1-8, then $T\textbf{-}GIC2=n^{-\#(M_{k})/n}\times GIC(M_{k})$ will consistently select the true model $M_{k_{0}}$.
@@ 4.1 $N\times t$ PDF
We conduct 100 replications to evaluate the performance of the MGICE, with sample size $n=1000,3000,\text{and }5000$, where the true parameter values are set to $(\mu^{*},s^{*},\alpha^{*},k^{*})=(0.3,0.5,0.5,1.5)$. Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution. For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm (Kingma and Ba, 2015) to jointly optimize all parameters. Similar to Baker (2022), for all experiments involving the $N\times t$ PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g. $\alpha=0.25$ or 1, and $k=1$ or 2. Table 4 presents the average MGICE and its SD, showing good overall consistency of MGICE, although there is still room for improvement for the parameters $s^{*},\alpha^{*},$ and $k^{*}$.
@@ 4.2 AR Model with $N\times t$ Noise
We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $N=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$. In each replication, the first 200 data are discarded to ensure stationarity.
@@ 4.3 Polynomial Regression Model with $N\times t$ Distributed Errors
We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $n=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
@@ 4.4 A Bivariate Model with a Von Mises PDF
We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using T-GIC1 and T-GIC2, with sample size $n=300,500,\text{and }1000$. The true parameter dimension is set to 5, with true parameter $\theta^{*}=(2.0,1.0,1.5,2.5,3.0)^{T}$. Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $[0,2\pi)\times[0,2\pi)$. For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Nocedal and Wright, 2006) to jointly optimize all parameters, with all parameters initialized to zero.
@@ 5.1 Finance Data
where $\varepsilon_{t}\sim N(0,\sigma^{2})$. Tables 10 and 11 summarize the results. Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model. However, under $N\times t$ noise, T-GIC1 selects a sixth-order AR model, while T-GIC2 selects a second-order AR model, both yielding estimates of $\alpha$ and $k$ similar to Baker (2022). Since AR models offer a convenient framework for prediction, Table 12 presents a comparison of out-of-sample performance between the two fitted models, either chosen by AIC vs BIC or by T-GIC1 vs T-GIC2. In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling $m$-step-ahead forecast. It is interesting to observe, for T-GIC1 vs T-GIC2, that the fitted AR(6) model performs better in forecasting for the short term than the fitted AR(2) model but worse for the longer term. However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly. Overall, Table 12 suggests that the effect of heavy-tailed innovation kicks in for the longer term forecasting.
@@
Table 12: MSE and the ratio in the rolling m-step-ahead forecast on AR models. (Row-wise minimum values are underlined.)
$m$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
@@ A.2 Proof of Theorem 1
First, we observe that for any model $M$ satisfying Assumptions 6-7 and assuming the validity of the law of large numbers, then it follows from routine analysis that $GIC(M(\theta))=P_{n}\{W(x,p_{M(\theta)})\}$ converges uniformly in probability to its population version $GIC_{\infty}(M(\theta))=P\{W(x,p_{M(\theta)})\}$. It follow from Assumption 7 that $GIC_{\infty}(M_{k}(\theta_{k}))$ is a Lipschitz-continuous, as a function of the parameter $\theta_{k}$, hence it attains its maximum value, denoted by $\mathcal{M}_{k}$, owing to the compact parameter space assumption (aka Assumption 6). Suppose $k_{0}$ is the smallest $k$ such that $M_{k}$ contains the true model. Then, $\mathcal{M}_{k}<\mathcal{M}_{k_{0}}$ for all $1\leq k<k_{0}$, whereas $\mathcal{M}_{k}=\mathcal{M}_{k_{0}}$ otherwise. Note that from Proposition 3 in Cheng and Tong (2024), $\mathcal{M}_{k_{0}}=H_{G}(p^{*})>0$ where $p^{*}$ is the true population pdf. Therefore, if for any $k$, $C(n,k)\to 1$ as $n\to\infty$, then the maximum T-GIC model selection criterion will not select any $k<k_{0}$, in probability.
Henceforth, consider the case that $k\geq k_{0}$. Since $\mathcal{M}_{k_{0}}>0$, $GIC(k)$ is positive, in probability, i.e., $GIC(k)>0$ holds with probability approaching 1 as sample size increases without bound. For ease of exposition, we shall assume that $GIC(k)$ is positive. Let $k>k_{0}$ be fixed. Consider the increment $D=\log\{T\text{-}GIC(k)\}-\log\{T\text{-}GIC(k_{0})\}=\log C(n,k)-\log C(n,k_{0})+\log GIC(k)-\log GIC(k_{0})$. By Proposition 1, $n\times\{\log GIC(k)-\log GIC(k_{0})\}$ converges weakly to some non-negative distribution. Consequently, $D=\log C(n,k)-\log C(n,k_{0})+O_{p}(1/n)$ so that if $n\times\log\{C(n,k)/C(n,k_{0})\}\to-\infty$ as $n\to\infty$, $D$ is negative in probability for $k>k_{0}$. This completes the proof of the consistency of the proposed T-GIC model selection criterion.